OpenStudy (anonymous):

how do i solve 17^{1-x}=7

OpenStudy (anonymous):

Take the log (any base) of each side, first.

OpenStudy (anonymous):

Use the law of logs to "bring down" the (1-x). Divide by log(17) Should be pretty straight forward to solve for x from there.

OpenStudy (anonymous):

change of base formula \[\large b^A=B\iff A=\frac{\log(A)}{\log(B)}\]

OpenStudy (anonymous):

\[\large 17^{1-x}=7\iff 1-x=\frac{\log(7)}{\log(17)}\] then algebra and a calculator

OpenStudy (anonymous):

Sorry, Im still not seeing it. Please explain